[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO biweekly status report

   Compton Observatory Science Report #174, Friday January 6, 1995
      Chris Shrader, Compton Observatory Science Support Center

         Questions or comments can be sent to the CGRO SSC.
          Phone:    301/286-8434
          e-mail:   NSI_DECnet: GROSSC::SHRADER 
          Internet: shrader@grossc.gsfc.nasa.gov

                         Instrument Reports


EGRET operations were normal this biweekly period. Delivery of data
to the GRO SSC remains on schedule. Interaction with guest
investigators continues at a good level.

Several papers on active galaxies have been accepted by the
Astrophysical Journal. Work on the long, complex analysis of the all
sky data to determine the diffuse radiation as a function of position
is nearing completion; preliminary reports have been presented at
scientific meetings, and papers on the galactic and presumably
extragalactic components should be submitted to scientific journals
in the near future.

The deep exposure to the Virgo region will be completed next Tuesday. 
The results of this study when it is finished should add
significantly to our understanding of active galactic nuclei.


In the past two weeks, OSSE detector #1 has failed to step twice, on
94/358 (24 Dec) and 95/002 (2 Jan).  There have now been three such
occurrences since launch; the first was 16 Dec 1994.  The motor drive
positioning process, which in normal operation moves the detectors
every 2 minutes, takes the detector off-line when it detects a
positioning error.  We restored the detector to normal operation
without any difficulty.  The circumstances of all three failures
suggest that it is not a mechanical problem with the drive or drive
motor, and we are investigating possible electronic or software

OSSE is currently in normal operations on all four detectors.

In viewing period 408 (3-10 Jan), the Z-axis targets are 3C 273 and
3C 279 (PI team), and the X-axis target is H1822-37.1 (Guest
Investigator E. Liang). When neither scheduled target is visible, the
targets are NGC 3227 and PKS 2155-304.

Three very interesting OSSE papers have been accepted for publication
recently, summarizing the long-term (spanning 18 months) spectral
behavior of 3C 273 (Johnson et al.), observations of the starburst
galaxies NGC 253 and M82 (Bhattacharya et al.), and the absence of a
red-shifted annihilation feature in the spectrum of 1E 1740.7-2942
(Jung et al.) on 20 Sep 1992.  With OSSE substantially greater
sensitivity, this last result is in direct conflict with the report
by SIGMA of transient emission on that day.

Data from viewing periods 303.2 and 303.4 were delivered to the
Compton GRO Science Support Center archive this week.  The targets
during these viewing periods were N Cyg 1992, NGC 5548, Cyg X-1, and


The COMPTEL instrument is performing well and continues routine
observations (and sends New Year's greetings to all its friends and

There will be several presentations by members of the COMPTEL
collaboration, and Guest Investigators, at next week's meeting of the
AAS in Tuscon.  Topics include recent spectral results for Cygnus
X-1, the current state of the BATSE/COMPTEL/NMSU burst rapid-response
network, and the latest results on searches for gamma-ray lines, and
signs of radiative neutrino decay, from supernovae.

We remind colleagues that abstract deadlines are approaching for
contributions to several conferences sponsored in part by the
collaboration: the 29th ESLAB Symposium, "Toward the Source of
Gamma-Ray Bursts" (meeting dates 25-27 April 1995; abstracts due
January 13); the Second Waterville Valley Workshop on High-Energy
Solar Phenomena (meeting dates 6-10 March 1995; abstracts due
February 1); and the Third Compton Syposium, in Munich (meeting dates
12-14 June 1995; abstracts due March 31).  Refer to the GRONEWS
on-line bulletin board for further details.


At upcoming AAS Meeting in Phoenix BATSE team scientists will be
presenting results from BATSE observations of the transient x-ray /
radio jet sources GRO J1655-40 and GRS 1915+105.  

The following sources were detected by the BATSE pulsed source
monitor in the past two weeks: 4U 1626-67, OAO 1657-415, GX 1+4, Vela
X-1, and GX 301-2. The x-ray binary GX 1+4 is still spinning up. The
outburst that began in October continues, with typical pulsed fluxes
of 80 mCrab (20-100 keV).

As of January 4th, BATSE has detected 1199 cosmic gamma-ray bursts
out of a total of 3239 on-board triggers in 1352 days of operation.
There have been 741 triggers due to solar flares with emission above
60 keV. Since the BATSE  burst trigger criteria were changed on
September 19th 1994, there have been 15 triggers due to terrestrial
gamma-ray flashes.